What is the equation of perpendicalur/paralel to given line through the point (3,2) and given line equation y = 5x-3.
Write down the equation in y = mx=b form.&
1
Expert's answer
2010-11-05T01:48:03-0400
Paralel lines have the same slope:
m1=m2 = 5.
for (3,2) : 2 = 5*3 + b; b = -13.
Thus the equation of paralel line is
y = 5x - 13.
The slope of perpendicular line equals to negative reciprocal of the slope of given line:
m2 = -1/5;
for (3,2): 2 = -1/5 * 3 + b; b = 13/5.
The equation of the perpendicular line is
y = -1/5 + 13/5.